# Feel-Good Thompson Sampling for Contextual Bandits: a Markov Chain Monte Carlo Showdown

### v1

@@ E.2 Logistic Bandits
Low-dimensional setting. For the $d=20$ setting, Tables 22 and 21 provide our experiments over logistic bandits for $\beta=10^{3}$ and $\beta=1$ (respectively), where we modify the feel-good parameter $\lambda$ in the loss-likelihood for values $\lambda\in\{0.01,0.1,0.5,1.0\}$. As before, we see that setting the feel-good parameter $\lambda$ to be 0.01 in FGMALATS and SFGMALATS outperforms the case where $\lambda=0$ (as well as other choices of $\lambda)$.
@@
High-dimensional setting. For the $d=40$ setting, Tables 24 and 23 provide our experiments over logistic bandits for $\beta=10^{3}$ and $\beta=1$ (respectively), where we modify the feel-good parameter $\lambda$ in the loss-likelihood for values $\lambda\in\{0.01,0.1,0.5,1.0\}$. Here, setting the feel-good parameter $\lambda$ to be either 0.01 or 0.1 favors each algorithm differently. These benefits are most evident in FGMALATS, SFGLMCTS, and SFGMALATS.
@@
FGLMCTS 1486.4 $\pm$ 114 1531.5 $\pm$ 128 1606.5 $\pm$ 120 1521.5 $\pm$ 153.4
FGMALATS 359.6 $\pm$ 58.2 372.1 $\pm$ 61.1 401 $\pm$ 56.6 428 $\pm$ 43.5
PFGLMCTS 2315 $\pm$ 207.2 2523.4 $\pm$ 222.3 2787.6 $\pm$ 348.2 2816.6 $\pm$ 281.8
PSFGLMCTS $3206.1\pm 200.1$ $1572.3\pm 100.1$ $1532\pm 230.7$ $1723.7\pm 139.9$
SFGLMCTS $475.3\pm 46.4$ $480.1\pm 52.8$ $527.7\pm 85.5$ $540\pm 91.1$
SFGMALATS $464.4\pm 78.2$ $399.1\pm 62.1$ $416\pm 68$ $385.5\pm 58.2$
@@ E.5 Neural Bandits
This subsection provides our main experimental results for our neural bandit experiments. Tables 30 and 31 detail the results of our experiments on classification tasks on 5 UCI datasets and the MNIST_784 dataset. Our experiments show that FGLMCTS loses its competitive edge on the neural tasks: for instance, the simple/cumulative regret of SFGLMCTS and FGLMCTS is much worse than vanilla LMCTS on a number of tasks.
@@ Appendix F Datasets
Table 36: Statistics of the benchmark datasets. Context dimension equals #$\text{\#arms}\times\text{\#attributes}#$.
@@
Wheel-2-$\delta$a 2 5 10 5000
Wheel-2-$\delta$b 2 5 10 5000
Wheel-2-$\delta$c 2 5 10 5000
Wheel-2-$\delta$d 2 5 10 5000
@@ F.1 Real-World Datasets
[…]
@@
Jester. Following Riquelme et al. [2018], we created a recommendation system bandit problem as follows. The Jester Dataset (Goldberg et al. [2001]) provides continuous ratings in $[-10,10]$ for $100$ jokes from a total of 73421 users. We find a complete subset of $T=19181$ users rating all $40$ jokes. As in Riquelme et al. [2018], we take $d=32$ of the ratings as the context of the user, and $k=8$ as the arms. The agent recommends one joke, and obtains the reward corresponding to the rating of the user for the selected joke.
@@
Telescope. The MAGIC Telescope dataset [Asuncion et al., 2007] comprises of $d=10$ real-valued features on cosmic ray events (length, energy, etc.), and a binary label for whether the event is a gamma-ray or a hadron. As in before, the agent obtains a reward $1$ for making the right prediction, and 00 otherwise. We set $T=10000$.
MNIST. MNIST [LeCun, 2010] comprises of $d=784$ real-valued features corresponding to flattened $26\times 26$ images of $k=10$ hand-drawn digits from 00 to $9$. The task again is to make a prediction $a\in\{0,\dots,9\}$ of the digit in the image, where the agent gets reward $1$ for making the right prediction and 00 otherwise. We set the horizon of the game to be $T=10000$.
@@ F.2 Synthetic Datasets
[…]
@@
[…]
@@
    Wheel-2-$\delta$a: set $\delta=0.01$ in the wheel bandit environment.
@@
    Wheel-2-$\delta$b: set $\delta=0.1$ in the wheel bandit environment.
@@
    Wheel-2-$\delta$c. set $\delta=0.5$ in the wheel bandit environment.
@@
    Wheel-2-$\delta$d. set $\delta=0.99$ in the wheel bandit environment.

### v3

@@ E.2 Logistic Bandits
Low-dimensional setting. For the $d=20$ setting, Tables 24 and 23 provide our experiments over logistic bandits for $\beta=10^{3}$ and $\beta=1$ (respectively), where we modify the feel-good parameter $\lambda$ in the loss-likelihood for values $\lambda\in\{0.01,0.1,0.5,1.0\}$. As before, we see that setting the feel-good parameter $\lambda$ to be 0.01 in FGMALATS and SFGMALATS outperforms the case where $\lambda=0$ (as well as other choices of $\lambda)$.
@@
High-dimensional setting. For the $d=40$ setting, Tables 26 and 25 provide our experiments over logistic bandits for $\beta=10^{3}$ and $\beta=1$ (respectively), where we modify the feel-good parameter $\lambda$ in the loss-likelihood for values $\lambda\in\{0.01,0.1,0.5,1.0\}$. Here, setting the feel-good parameter $\lambda$ to be either 0.01 or 0.1 favors each algorithm differently. These benefits are most evident in FGMALATS, SFGLMCTS, and SFGMALATS.
@@
FGLMCTS $1486.4\pm 114.0$ $1531.5\pm 128.0$ $1606.5\pm 120.0$ $1521.5\pm 153.4$
FGMALATS 359.6 $\pm$ 58.2 $372.1\pm 61.1$ $401.0\pm 56.6$ $428.0\pm 43.5$
PFGLMCTS $2315.0\pm 207.2$ $2523.4\pm 222.3$ $2787.6\pm 348.2$ $2816.6\pm 281.8$
PSFGLMCTS $3206.1\pm 200.1$ $1572.3\pm 100.1$ $1532.0\pm 230.7$ $1723.7\pm 139.9$
SFGLMCTS $475.3\pm 46.4$ $480.1\pm 52.8$ $527.7\pm 85.5$ $540.0\pm 91.1$
SFGMALATS $464.4\pm 78.2$ 399.1 $\pm$ 62.1 416.0 $\pm$ 68.0 385.5 $\pm$ 58.2
@@ E.5 Neural Bandits
This subsection provides our main experimental results for our neural bandit experiments. Tables 32 and 33 detail the results of our experiments on classification tasks on 7 UCI datasets ((Adult, Shuttle, MagicTelescope, Mushroom, Covertype), the RestaurantRatings (SCI) and Jester datasets), Financial dataset and two vision benchmarks, MNIST_784 and CIFAR-10. Our experiments show that FGLMCTS generally loses its competitive edge on the neural tasks: for instance, the simple/cumulative regret of SFGLMCTS and FGLMCTS is much worse than vanilla LMCTS on a number of tasks.
@@ Appendix F Datasets
Table 38: Statistics of the benchmark datasets. Context dimension equals $\text{\#arms}\times\text{\#attributes}$.
@@
[…]
@@ F.1 Real-World Datasets
Mushroom. The Mushroom Dataset [Lincoff, 1997] contains 22 mushroom features and labels for whether the mushroom is safe to consume: $\{$“poisonous”, “safe” }\}. As in Blundell et al. [2015], we create a bandit problem where the agent must decide whether to eat the mushroom. In this setting, eating a safe mushroom delivers a deterministic reward of +5+5, and eating a poisonous mushroom provides a randomized reward of +5+5 with probability $1/2$ and 35-35 reward with probability $1/2$. If the agent does not eat a mushroom, then the reward is 0. In this case, we set $T=10000$.
@@
Jester. Following Riquelme et al. [2018], we created a recommendation system bandit problem as follows. The Jester Dataset (Goldberg et al. [2001]) provides continuous ratings in [10,10][-10,10] for $100$ jokes from a total of 73421 users. We find a complete subset of $T=19181$ users rating all $40$ jokes. As in Riquelme et al. [2018], we take $d=32$ of the ratings as the context of the user, and $k=8$ as the arms. The agent recommends one joke, and obtains the reward corresponding to the rating of the user for the selected joke.
@@
Telescope. The MAGIC Telescope dataset [Asuncion et al., 2007] comprises of $d=10$ real-valued features on cosmic ray events (length, energy, etc.), and a binary label for whether the event is a gamma-ray or a hadron. As in before, the agent obtains a reward $1$ for making the right prediction, and 0 otherwise. We set $T=10000$.
MNIST_784. MNIST [LeCun, 2010] comprises of $d=784$ real-valued features corresponding to flattened $26\times 26$ images of $k=10$ hand-drawn digits from 0 to $9$. The task again is to make a prediction $a\in\{0,\dots,9\}$ of the digit in the image, where the agent gets reward $1$ for making the right prediction and 0 otherwise. We set the horizon of the game to be $T=10000$.
CIFAR-10. CIFAR-10 [Krizhevsky, 2009] comprises of $d=3072$ real-valued features corresponding to flattened $32\times 32\times 3$ RGB images of $k=10$ object classes (airplanes, cars, birds, cats, deer, dogs, frogs, horses, ships, and trucks). The task is to make a prediction $a\in\{0,\dots,9\}$ of the object class in the image, where the agent gets reward $1$ for making the correct prediction and 0 otherwise. We set the horizon of the game to be $T=25000$.
RestaurantRatings. Restaurant [Medelln and Serna, 2011] contains of $d=128$ real-valued features corresponding to user and restaurant contextual information from the UCI Restaurant and Consumer Data. The dataset contains $k=127$ restaurants and the task is to make a recommendation $a\in\{0,\dots,126\}$ for a restaurant, where the agent gets reward $1$ for a positive rating and 0 otherwise. We set $T=10000$.
@@ F.2 Synthetic Datasets
Linear Contextual Bandit. Fix a dimension $d$ and horizon T=10  [^0] T=10\,000. The linear contextual bandit environment is given as follows: at round $t\in[T]$, the agent observes a context $\mathcal{X}_{t}\sim\mathcal{N}(\mathbf{0}_{4},\mathbf{I}_{4})$, chooses action $x_{t}\in[K]$ with $K=5$, and receives a noisy linear reward $r_{t}=\phi(\mathcal{X}_{t},x_{t})^{\top}\theta^{\star}+\varepsilon_{t}$}, where $\varepsilon_{t}\sim\mathcal{N}(0,\sigma^{2}),\sigma=0.5$} and $\theta^{\star}\in\mathbb{R}^{d}$}. We place a Gaussian prior, $\theta_{0}\sim\mathcal{N}\bigl(\mathbf{0}_{d},\sigma_{0}^{2}\mathbf{I}\bigr)$} with $\sigma_{0}=0.01$. The feature map $\phi:\mathbb{R}^{4}\times[K]\to\mathbb{R}^{d}$ is the standard block concatenation $(\phi(\mathcal{X}_{t},0),\dots,\phi(\mathcal{X}_{t},K))$ where $\phi(\mathcal{X}_{t},i)=e_{i}\cdot\mathcal{X}_{t}$} (where $e_{i}$ is the $i$’th standard basis vector). Now, $\theta^{\star}$ naturally decomposes into $k$ context-specific blocks. Since the likelihood and prior are Gaussian, TS admits a closed-form posterior update, giving a convenient ground-truth baseline for our approximate sampling methods. We derive the following synthetic datasets:
@@
Logistic Contextual Bandit. Fix a dimension $d$. We further consider a logistic contextual bandit to introduce non-linear reward dependencies. With T=10 000{T=10\,000} and K=50{K=50} arms, at each round $t\in[T]$}, the learner observes a collection of arm‐specific context vectors $\mathcal{X}_{t,a}\sim\mathcal{N}\bigl(\mathbf{0}_{d},\,I_{d}\bigr),$} where $a=1,\dots,50$} each of which is then normalized to unit norm. The learner selects an arm $a_{t}\in[K]$ and obtains a Bernoulli reward $r_{t}\sim\mathrm{Bern}\bigl(\sigma\bigl(\phi(\widetilde{\mathcal{X}}_{t,a_{t}})^{\top}\theta^{\star}\bigr)\bigr),$} where $\sigma(u)=\frac{1}{1+e^{-u}}$}, $\theta^{\star}\sim\mathcal{N}\bigl(\mathbf{0}_{d},\mathbf{I}_{d}\bigr)$} (scaled to unit norm), and $\smash{\phi:\mathbb{R}^{d}\to\mathbb{R}^{d}}$ where $\smash{\phi(\mathcal{X}_{t})=\mathcal{X}_{t}}$ is the identity feature map extracting the $d$-dimensional context for each arm. We place a Gaussian prior $\theta_{0}\sim\mathcal{N}\bigl(\mathbf{0}_{d},\sigma_{0}^{2}\mathbf{I}\bigr)$}, with $\sigma_{0}=0.01$. Similarly, we derive the following synthetic datasets:
@@
[…]
@@
[…]
@@
[…]
@@
[…]